# HR-INR: Continuous Space-Time Video Super-Resolution via Event Camera

### v1

@@ Abstract
Continuous space-time video super-resolution (C-STVSR) aims to simultaneously enhance video resolution and frame rate at an arbitrary scale. Recently, implicit neural representation (INR) has been applied to video restoration, representing videos as implicit fields that can be decoded at an arbitrary scale. However, the highly ill-posed nature of C-STVSR limits the effectiveness of current INR-based methods: they assume linear motion between frames and use interpolation or feature warping to generate features at arbitrary spatiotemporal positions with two consecutive frames. This restrains C-STVSR from capturing rapid and nonlinear motion and long-term dependencies (involving more than two frames) in complex dynamic scenes. In this paper, we propose a novel C-STVSR framework, called HR-INR, which captures both holistic dependencies and regional motions based on INR. It is assisted by an event camera – a novel sensor renowned for its high temporal resolution and low latency. To fully utilize the rich temporal information from events, we design a feature extraction consisting of (1) a regional event feature extractor – taking events as inputs via the proposed event temporal pyramid representation to capture the regional nonlinear motion and (2) a holistic event-frame feature extractor for long-term dependence and continuity motion. We then propose a novel INR-based decoder with spatiotemporal embeddings to capture long-term dependencies with a larger temporal perception field. We validate the effectiveness and generalization of our method on four datasets (both simulated and real data), showing the superiority of our method.
@@ 1 Introduction
The real world’s visual information, e.g., edge and object motion, is continuous, spanning both time and space dimensions. However, the limited I/O bandwidth and sensor size of modern systems [13, 49, 47] confines us to record videos at low frame rates and fixed resolutions. This limitation has profound repercussions across various computer vision applications, e.g., encompassing immersive experiences in virtual reality [77, 31], traffic analysis in autonomous driving [84, 81]. To address this limitation, recent research works [10, 8] have explored restoring videos with continuous resolutions and frame rates, referred to as Continuous Space-Time Video Super-Resolution (C-STVSR).
Recently, implicit neural representation (INR) has been applied to video restoration: it represents images or videos as neural fields that can be decoded at any resolution with a pointwise MLP decoder [9, 74, 38, 4, 74, 7]. One of the seminal INR approaches is LIIF [9], which is designed for arbitrary-scale image SR. This line of research soon extended to the video domain. In this context of C-STVSR, VideoINR [10] employs a fixed STVSR model [69] that extracts features from two consecutive video frames. Then, it introduces a temporal INR to generate inverse backward warping optical flow [42] to warp features. Lastly, it employs a spatial INR, similar to LIIF, to decode the RGB frame with arbitrary resolution. Building upon this, MoTIF [8] improves VideoINR by using forward motion estimation, reducing gaps and holes in the temporal INR, which are typically caused by the randomness and discontinuities associated with backward warping [22, 46]. These methods depend solely on two successive RGB frames, rendering the task of predicting inter-frame motions ill-posed. Consequently,, it becomes challenging to accurately capture highly dynamic motion (e.g., regional high-speed or nonlinear movements) and to model long-term dependencies that extend across more than four frames.
@@
Figure 1: With event data as guidance, our method (HR-INR) takes in videos with low frame rates and resolution (a) and produces continuous space-time videos with arbitrary frame rate and resolution (b). As shown in (c), our method is able to recover the rotation of the bicycle wheels, which is unachievable by the prior method VideoINR [10].
Motivation and Contributions. Event cameras are bio-inspired sensors, known for their high temporal resolution and low latency ($<1us$[83, 14, 63]. Recent research has demonstrated the potential of events in guiding various video super-resolution (VSR) [38, 23] and video frame interpolation (VFI) tasks [62, 61, 20, 75]. However, utilizing events to facilitate joint video super-resolution and frame interpolation is a challenging area yet to be explored.
This paper introduces HR-INR, a novel INR-based method that leverages events for jointly guiding VSR and VFI. It adeptly captures regional, rapid motion and holistic, long-range motion dependencies, as shown in Fig. 1. To capture the regional motion, we propose Temporal Pyramid Representation (TPR) to construct a time-series pyramid structure around the pivotal timestamp of events (Sec. 3.1). Different from the evenly divided timeline representations, like voxel grids [62, 61, 14], time surfaces [53], time moments [17, 38] and symmetric cumulative [57], TPR offers finer temporal granularity with less complexity and effectively captures rapid motion changes, as shown in Fig. 2 (a). Additionally, our method is capable of processing multiple frames and their associated events, which empowers it to estimate holistic, long-range motion that involves more than two frames.
@@
Accordingly, we design two specialized feature extractors: the regional event feature extractor (RE) and the holistic event-frame feature extractor (HE), see Sec. 3.2. Both extractors are grounded in the Swin-Transformer architecture [32, 36, 33], renowned for its efficacy and efficiency in video enhancement tasks. RE is a lightweight network designed specifically for extracting local information from our event TPR. Meanwhile, HE employs a more sophisticated approach, utilizing long-term and multi-scale fusion strategies to integrate both events and frames. Consequently, our training and inference strategy requires HE to be invoked only once for multi-frame interpolation. Subsequently, the extracted features from RE and HE are fused as the output of the feature extraction module.
@@
After fusing the regional and holistic features, we propose a novel INR-based spatial-temporal decoding module (Sec. 3.3). Our motivation is to avoid gaps and holes typically found in optical flow-based warping and multi-frame fusion, as identified in previous research [8, 10, 20, 61]. To accomplish this, we propose an implicit temporal embedding designed to transform timestamps into focused attention vectors on long-term features. This approach ensures attention is also given to long-distance dependencies, which is crucial for effectively modeling long-term temporal dependencies. Subsequently, inspired by LIIF [9], we employ spatial embedding to achieve arbitrary up-sampling in the spatial dimension.
@@
We conducted experiments on two simulated and two real-world datasets. The results validate the superiority of our method and its excellent generalization capabilities on real-world datasets. Our approach is the first event-based method to achieve continuous space-time video super-resolution, surpassing frame-based methods, as shown in Fig. 1. It also excels in individual VSR and VFI metrics compared to previous event-based methods.
@@ 2 Related Works
Space-time Video Super-Resolution aims to enhance the resolution and frame rate of a video simultaneously [18, 27, 69, 70]. In comparison to two-stage solutions, where VFI [22, 72, 42, 43, 11] and VSR [34, 73, 76, 65, 60, 21] methods are applied sequentially, simultaneous space-time video super-resolution reduces cumulative errors and leverages the natural relations between VFI and VSR methods. Zooming Slow-Mo [69] uses temporal interpolation to generate missing frames and aligns temporal information using a deformable ConvLSTM network. Similarly, TMNet [70] extracts short-term and long-term motion cues in videos by modulating convolution kernels. However, these methods cannot simultaneously achieve spatiotemporal resolution across arbitrary scales.
INR for VFI and VSR have achieved space-time video super-resolution with arbitrary resolutions [10, 8] by learning videos implicit neural representations (INRs). These methods primarily estimate optical flows from two consecutive frames to warp features into arbitrary space-time coordinates, which are then decoded using MLP layers. However, by relying on just two consecutive frames, these methods cannot inherently model long-term motions (involving three or more frames) and fail to accurately capture local, inter-frame, non-linear motions due to missing inter-frame information.
Event-guided VFI and VSR seek to boost performance by incorporating the biologically inspired event cameras [83]. Previous works have demonstrated the potential of event-guided VFI, which mainly focus on modeling non-linear motion with events [44, 62, 68, 61, 20, 55, 54]. EFI [44] exclusively adopts the synthesis approach for intermediate frame generation. TimeLens [62] and TimeLens++ [61] employ events to model nonlinear motion correlations, integrating both synthesis and warping-centric approaches. Building on these advancements, CBM-Net [28], introduces a motion field to handle complex movements. However, while these VFI methods utilize events to capture inter-frame motion, they fail to establish long-term dependencies beyond two frames and support simultaneous VSR. The realm of event-guided video super-resolution has also been explored. E-VSR [23] highlights that high-frequency temporal information from events is beneficial to recovering high-frequency spatial information. Like our work, EG-VSR [38] employs events to comprehend INR, allowing for video ups-sampling with arbitrary scale. However, the INR of EG-VSR cannot interpolate frames. Contrasting these methods, we pioneer using events to enable concurrent VSR and VFI across arbitrary spatial-temporal scales, i.e.,, C-STVSR.
Video Long-term Dependence Modeling is a crucial aspect of VSR and VFI. For instance, BasicVSR [5] and BasicVSR++[6] enhance VSR performance by processing multi-frame inputs through an alignment module to model long-term motion correlations. Similarly, in the VFI domain, many methods [58, 41, 80] employ RNN or LSTM to model sequences of frames, capturing long-term dependencies effectively. Furthermore, Zooming Slow-Mo [69], TMNet [70], and RSST [33] leverage multi-frame inputs in the joint task of VSR and VFI, showcasing the importance of integrating multiple frames for improved modeling of video dynamics. However, current C-STVSR methods [10, 8], and event-based VFI methods [62, 20, 61, 28], primarily rely on estimating optical flow between two consecutive frames. Therefore, they are challenging to handle multi-frames as inputs, inherently undermining their capability to model long-term dependencies.
@@ 3 Proposed Framework
Our proposed HR-INR framework is depicted in Fig. 2. The inputs of this framework are RGB frames $\boldsymbol{I}_{in}=\{I_{in}\}_{i}^{N_{in}}\in R^{N_{in}\times H\times W\times 3}$ and associated events $\boldsymbol{E}$. $H$ and $W$ denote the spatial resolution of frames and events. $3$ means three channels of RGB. $N_{in}$ denotes the input number of frames. Furthermore, the framework outputs a video with an arbitrary frame rate and spatial resolution. In particular, we consider the output video as $\boldsymbol{I}_{out}=\{I_{out}\}_{i}^{N_{out}}$, consists of $N_{out}$ frames, each with a resolution of $(s\times H)\times(s\times W)$, where $s$ represents the up-sampling scale greater than $1$. For the output $N_{out}$ frames, we denote the time corresponding to each frame as $\boldsymbol{T}=\{t\}_{i}^{N_{out}}$. For convenience, we also record the up-sampling scale $s$ and the time $\boldsymbol{T}$ as a part of inputs. Therefore, the mapping function $f_{hr}(.)$ of C-STVSR can be described by Eq. 1.
$\boldsymbol{I}_{out}=f_{hr}\left(\boldsymbol{I}_{in},\boldsymbol{E},s,\boldsymbol{T}\right)$ (1)
@@
Input Frames and Events: Our input frames, $\boldsymbol{I}_{in}$, consist of multiple frames with timestamps normalized to the $[0,1]$ interval. We consider the events, $\boldsymbol{E}$, occurring within this time range. Each event point can be represented by $(x,y,t,p)$, signifying a change in pixel intensity at coordinates $(x,y)$ at time $t$; here, $p$ is $+1$ for increased brightness and $-1$ for a decrease. Benefiting from the event camera’s high temporal resolution and low latency ($<1us$), these event points are effectively considered continuous along the timeline. For a detailed exposition of the principles of event generation, please refer to the Suppl. Mat..

### v2

@@ Abstract
Continuous space-time video super-resolution (C-STVSR) aims to simultaneously enhance video resolution and frame rate at an arbitrary scale. Recently, implicit neural representation (INR) has been applied to video restoration, representing videos as implicit fields that can be decoded at an arbitrary scale. However, existing INR-based C-STVSR methods typically rely on only two frames as input, leading to insufficient inter-frame motion information. Consequently, they struggle to capture fast, complex motion and long-term dependencies (spanning more than three frames), hindering their performance in dynamic scenes. In this paper, we propose a novel C-STVSR framework, named HR-INR, which captures both holistic dependencies and regional motions based on INR. It is assisted by an event camera – a novel sensor renowned for its high temporal resolution and low latency. To fully utilize the rich temporal information from events, we design a feature extraction consisting of (1) a regional event feature extractor – taking events as inputs via the proposed event temporal pyramid representation to capture the regional nonlinear motion and (2) a holistic event-frame feature extractor for long-term dependence and continuity motion. We then propose a novel INR-based decoder with spatiotemporal embeddings to capture long-term dependencies with a larger temporal perception field. We validate the effectiveness and generalization of our method on four datasets (both simulated and real data), showing the superiority of our method. The project page is available at https://github.com/yunfanLu/HR-INR.
journal: International Journal of Computer Vision
@@ 1 Introduction
Figure 1: With event data as guidance, our method (HR-INR) takes in videos with low frame rates and resolution (a) and produces continuous space-time videos with arbitrary frame rate and resolution (b). Demonstrating effective modeling of local nonlinear motion, our method uniquely showcases this with the bicycle example in (c), a feat unachievable by the prior method - VideoINR (Chen et al, 2022). As shown in (c), our method is able to recover the rotation of the bicycle wheels, which is unachievable by the prior method VideoINR (Chen et al, 2022).
The real world’s visual information, e.g., edge and object motion, is continuous, spanning both time and space dimensions. However, the limited I/O bandwidth and sensor size of modern systems (Delbracio et al, 2021; Parker, 2010) confines us to record videos at low frame rates and fixed resolutions. This limitation has profound repercussions across various computer vision applications, e.g., encompassing immersive experiences in virtual reality (Zhang, 2020; Lee et al, 2020), traffic analysis in autonomous driving (Zou et al, 2023; Zhao et al, 2019). To address this limitation, recent research works, e.g., VideoINR (Chen et al, 2022, 2023b), have explored restoring videos by implicit neural representation (INR), with continuous resolutions and frame rates, referred to as Continuous Space-Time Video Super-Resolution (C-STVSR).
In practice, INR-based methods represent images or videos as neural fields that can be decoded at any resolution using a pointwise MLP decoder (Cao et al, 2023; Chen et al, 2023a). Once an image or video is encoded into this neural field, it effectively becomes a continuous function over space and time. Consequently, querying any previously unseen coordinate $(x,y,t)$ yields the corresponding RGB value, thus enabling arbitrary-scale spatial upsampling and arbitrary-frame-rate temporal interpolation. One of the seminal INR approaches is LIIF (Chen et al, 2021), which is designed for arbitrary-scale image SR. This line of research soon extended to the video domain. In this context of C-STVSR, VideoINR (Chen et al, 2022) employs a fixed STVSR model (Xiang et al, 2020) that extracts features from two consecutive video frames. Then, it introduces a temporal INR to generate inverse backward warping optical flow (Niklaus and Liu, 2020) to warp features. Lastly, it employs a spatial INR, similar to LIIF, to decode the RGB frame with arbitrary resolution. Building upon this, MoTIF (Chen et al, 2023b) improves VideoINR by using forward motion estimation, reducing gaps and holes in the temporal INR, which are typically caused by the randomness and discontinuities associated with backward warping (Park et al, 2021). However, these methods depend solely on two successive RGB frames, rendering the task of predicting inter-frame motions ill-posed, as shown in Fig. 1 (c). Consequently, it becomes challenging to accurately capture highly dynamic motion (e.g., regional high-speed or complex nonlinear movements) and to model long-term dependencies that extend across more than four frames.
Motivation and Contributions: Event cameras are bioinspired sensors, known for their high temporal resolution and low latency (<1us<1us) (Gallego et al, 2020; Zheng et al, 2023; Lu et al, 2025b). Recent research has demonstrated the potential of events in guiding various video super-resolution (VSR) (Lu et al, 2023; Jing et al, 2021) and video frame interpolation (VFI) tasks (Tulyakov et al, 2021, 2022; He et al, 2022; Chen et al, 2025; Wang et al, 2025; Yan et al, 2025; Wei et al, 2025). Specifically, Jing et al (2021) demonstrated that the high temporal resolution of events can be converted into high spatial resolution of frames. Furthermore, Lu et al (2023) then extended this advantage to arbitrary-scale spatial upsampling. In the VFI, event streams provide motion information between frames. Tulyakov et al (2021, 2022) proposed dual-stream methods to leverage the temporal information from events for interpolation at any given moment. Kim et al (2023) modeled finer motions through motion field estimation. These methods have achieved significant progress in VFI and VSR tasks, and demonstrated how events can enhance both temporal information and spatial information. However, utilizing events to facilitate joint video super-resolution and frame interpolation remains a challenging area that has not been thoroughly explored. To end this, this paper introduces HR-INR, a novel INR-based method that leverages events for jointly guiding VSR and VFI through holistic and regional motion modeling. It adeptly achieves C-STVSR by capturing regional, rapid motion and holistic, long-range motion dependencies, as shown in Fig. 1. To capture the regional motion, we propose Temporal Pyramid Representation (TPR) to construct a time-series pyramid structure around the pivotal timestamp of events (Sec. 3.1). Different from the evenly divided timeline representations, like voxel grids (Tulyakov et al, 2021, 2022; Gallego et al, 2020), time surfaces (Sironi et al, 2018), time moments (Han et al, 2021; Lu et al, 2023) and symmetric cumulative (Sun et al, 2022), TPR offers finer temporal granularity with less complexity and effectively captures rapid motion changes, as shown in Fig. 2 (a). Additionally, to estimate holistic, long-range motion (involving more than two frames), our model can employ more frames (e.g., four frames) and their corresponding events. This makes it possible to capture the extended-duration motion and improves the accuracy and continuity of motion modeling.
@@
Accordingly, we design two specialized feature extractors: the regional event feature extractor (RE) and the holistic event-frame feature extractor (HE), see Sec. 3.2. Both extractors are grounded in the Swin-Transformer architecture (Liang et al, 2021; Liu et al, 2022a), renowned for its efficacy and efficiency in video enhancement tasks. The regional event feature extractor is a lightweight network designed specifically for extracting local information from our event TPR. Meanwhile, holistic event-frame feature extractor employs a more sophisticated approach, utilizing long-term and multi-scale fusion strategies to integrate both events and frames. Consequently, our training and testing strategy requires HE to be invoked only once for multi-frame interpolation. Subsequently, the extracted features from RE and HE are fused as the output of the feature extraction module.
@@
After fusing the regional and holistic features, we propose a novel INR-based spatial-temporal decoding module (Sec. 3.3). Our motivation is to reduce gaps and holes typically found in optical flow-based warping and multi-frame fusion, as show in Fig.  1 (c), which are also identified in previous research (Chen et al, 2023b, 2022; He et al, 2022; Tulyakov et al, 2022). To accomplish this, we propose an implicit temporal embedding designed to transform timestamps into focused attention vectors on long-term features. This approach is crucial for effectively modeling long-term temporal dependencies and capturing regional dynamic motion and edges, ensuring attention is also given to long-distance dependencies. Subsequently, we employ spatial embedding (Chen et al, 2021) to achieve arbitrary up-sampling in the spatial dimension.
@@
We conducted experiments on two simulated (Adobe240 (Su et al, 2017) and GoPro (Nah et al, 2017)) and two real-world datasets (BS-ERGB (Tulyakov et al, 2022) and CED (Scheerlinck et al, 2019a)). The results validate the superiority of our method and its excellent generalization capabilities on real-world datasets. Our approach is the first event-based method to achieve continuous space-time video super resolution, surpassing frame-based methods, as shown in Fig. 1. It also excels in individual VSR and VFI metrics compared to previous event-based methods.
@@ 2 Related Works
Since our research focuses on leveraging event data to guide the C-STVSR task, we categorize related works into five areas: (1) Frame-based VFI and VSR; (2) Frame-based Space-Time Video Super-Resolution; (3) INR for VFI and VSR; (4) Event-guided VFI and VSR; (5) Long-term Video Dependency Modeling. Sections (1) and (2) discuss the progress and limitations of frame-based methods, (3) highlights the role of INR in arbitrary scales, (4) covers event applications in VFI and VSR, and (5) emphasizes the importance of modeling long-term dependencies in videos.
(1) Frame-based VFI and VSR VFI generates intermediate frames, while VSR improves video spatial resolution. Early frame-based works addressed these tasks separately. VSR focuses on modeling motion relations between multiple frames for feature alignment and resolution enhancement. Early methods Bao et al (2019b); Wang et al (2019) used motion compensation and deformable convolutions, while later works (Chan et al, 2021, 2022a) explored longer temporal modeling. VFI, on the other hand, models inter-frame motion to generate intermediate frames, with early approaches relying on warping, kernels, or optical flow (Meyer et al, 2018; Xu et al, 2019; Dutta et al, 2021; Zhang et al, 2022). Despite attempts to model nonlinear motion, these methods remain limited by the absence of inter-frame data. Recent work, such as RSST (Liang et al, 2022), improves VFI by using multi-frame inputs. In summary, both VSR and VFI rely on inter-frame motion modeling but are constrained by missing inter-frame motion information, making it challenging to capture complex nonlinear motion.
@@
(2) Frame-based Space-time Video Super-Resolution aims to enhance the resolution and frame rate of a video simultaneously (Haris et al, 2020; Kim et al, 2020; Xiang et al, 2020; Xu et al, 2021). In comparison to two-stage solutions, where VFI  (Jiang et al, 2018; Xue et al, 2019; Niklaus and Liu, 2020; Niklaus et al, 2017; Cheng and Chen, 2020) and VSR  (Liu et al, 2018; Yang et al, 2021; Yue et al, 2022; Wang et al, 2021; Tian et al, 2020; Isobe et al, 2020) methods are applied sequentially, simultaneous space-time video super-resolution reduces cumulative errors and leverages the natural relations between VFI and VSR methods. Zooming Slow-Mo (Xiang et al, 2020) uses temporal interpolation to generate missing frames and aligns temporal information using a deformable ConvLSTM network. Similarly, TMNet (Xu et al, 2021) extracts short-term and long-term motion cues in videos by modulating convolution kernels. However, these methods cannot simultaneously achieve spatiotemporal resolution across arbitrary scales. i.e.,C-STVSR.
@@
(3) INR for VFI and VSR have achieved space-time video super-resolution with arbitrary resolutions (Chen et al, 2022, 2023b) by learning videos implicit neural representations (INRs). These methods primarily estimate optical flows from two consecutive frames to warp features into arbitrary space-time coordinates, which are then decoded using MLP layers. However, by relying on just two consecutive frames, these methods cannot inherently model long-term motions (involving three or more frames) and fail to accurately capture local, inter-frame, non-linear motions due to missing inter-frame information.
(4) Event-guided VFI and VSR seek to boost performance by incorporating the biologically inspired event cameras (Zheng et al, 2023; Lu et al, 2025a). Previous works have demonstrated the potential of event-guided VFI, which mainly focus on modeling non-linear motion with events  (Tulyakov et al, 2021, 2022; He et al, 2022; Song et al, 2023; Chen et al, 2025; Wang et al, 2025; Yan et al, 2025; Wei et al, 2025). EFI (Paikin et al, 2021) exclusively adopts the synthesis approach for intermediate frame generation. TimeLens (Tulyakov et al, 2021) and TimeLens++ (Tulyakov et al, 2022) employ events to model nonlinear motion correlations, integrating both synthesis and warping-centric approaches. Building on these advancements, CBM-Net (Kim et al, 2023), introduces a motion field to handle complex movements. E-CIR (Song et al, 2022) and DeblurSR (Song et al, 2023) use events to learn a function to map time to gray values, which can speed up the prediction time of each frame. However, while these VFI methods utilize events to capture inter-frame motion, they fail to establish long-term dependencies beyond two frames and support simultaneous VSR. The realm of event-guided video super-resolution has also been explored. E-VSR (Jing et al, 2021) highlights that high-frequency temporal information from events is beneficial to recovering high-frequency spatial information. EG-VSR (Lu et al, 2023) employs events to comprehend INR, allowing for video ups-sampling with arbitrary scale. Contrasting these methods, we pioneer using events to enable concurrent VSR and VFI across arbitrary spatial-temporal scales, i.e.,C-STVSR.
(5) Video Long-term Dependence Modeling is a crucial aspect of VSR and VFI. For instance, BasicVSR (Chan et al, 2021) and BasicVSR++(Chan et al, 2022b) enhance VSR performance by processing multi-frame inputs through an alignment module to model long-term motion correlations. Similarly, in the VFI domain, many methods (Suzuki and Ikehara, 2020; Nah et al, 2019; Zhang et al, 2020) employ RNN or LSTM to model sequences of frames, capturing long-term dependencies effectively. Furthermore, Zooming Slow-Mo (Xiang et al, 2020), TMNet (Xu et al, 2021), and RSST (Liang et al, 2022) leverage multi-frame inputs in the joint task of VSR and VFI, showcasing the importance of integrating multiple frames for improved modeling of video dynamics. However, current C-STVSR methods (Chen et al, 2022, 2023b), and event-based VFI methods (Tulyakov et al, 2021; He et al, 2022; Tulyakov et al, 2022; Kim et al, 2023), primarily rely on estimating optical flow between two consecutive frames, which is then used to warp these two frames to generate the intermediate frames. Therefore, they are challenging to handle multi-frames as inputs, inherently undermining their capability to model long-term dependencies.
@@ 3 Proposed Framework
Preliminary: Events are discrete points that capture the positive and negative changes in pixel brightness. Their generation depends on brightness alterations within the logarithmic domain. Specifically, an event point $e=(x,y,t,p)$ is triggered and logged upon meeting certain criteria. Suppose $L(x,y,t)$ represents the brightness at point $(x,y)$ at any given time $t$. The event is recorded if the absolute difference $\Delta L=log\left(L(x,y,t)\right)-L\left(x,y,t-\Delta t\right)$ surpasses a predetermined threshold $C$, as formulated as Eq. 1.
$L(x,y,t)=I(x,y,t^{\prime})\times\exp({C\int_{t^{\prime}}^{t}p~dt})$ (2)
Utilizing Eq. 2 along with corresponding events, intensity frames at a given time can be used to compute frames at alternate times, facilitating video frame interpolation. However, events are often noisy, and directly applying Eq. 2 may not yield optimal results (Pan et al, 2019). Therefore, employing neural networks for event-based frame interpolation has garnered significant attention (Paikin et al, 2021; Tulyakov et al, 2021; Zhang and Yu, 2022; Kim et al, 2023). Additionally, the high temporal resolution of events helps in converting to high spatial resolution, as demonstrated by previous studies (Lu et al, 2023; Jing et al, 2021). In conclusion, event signals play a crucial role in frame interpolation and video super-resolution tasks, making them a natural guide for continuous space-time video super-resolution.
Overview of HR-INR: Our proposed HR-INR framework is depicted in Fig. 2. The inputs of this framework are RGB frames $\bm{I}_{in}=\{I_{in}\}_{i}^{N_{in}}\in R^{N_{in}\times H\times W\times 3}$ and associated events $\bm{E}$. $H$ and $W$ denote the spatial resolution of frames and events. $3$ means three channels of RGB. $N_{in}$ denotes the input number of frames. Furthermore, the framework outputs a video with an arbitrary frame rate and spatial resolution. In particular, we consider the output video as $\bm{I}_{out}=\{I_{out}\}_{i}^{N_{out}}$, consists of $N_{out}$ frames, each with a resolution of $(s\times H)\times(s\times W)$, where $s$ represents the up-sampling scale greater than $1$. For the output $N_{out}$ frames, we denote the time corresponding to each frame as $\bm{T}=\{t\}_{i}^{N_{out}}$. For convenience, we also record the up-sampling scale $s$ and the time $\bm{T}$ as a part of inputs. Therefore, the mapping function $f_{hr}(.)$ of C-STVSR can be described by Eq. 3.
$\bm{I}_{out}=f_{hr}\left(\bm{I}_{in},\bm{E},s,\bm{T}\right)$ (3)
@@
Figure 3: Visualization of Event TPR across different time resolutions. The TPR is divided into 7 layers ($L=7$), with each layer having a time resolution that is $1/3$ of the previous layer ($r=3$). The time resolution for the first layer ($L_{0}$) is approximately $2\times\Delta t=1/9s$, and the time resolution for the seventh layer ($L_{6}$) is approximately $1/6561s$. Each layer visualizes the corresponding event data at a specific time resolution, demonstrating the effect of varying temporal resolution on event-based data.